Please? I need help? Thank you.
solve by graphing: x - 2y = 8
x + y = - 1

solve by substitution: 5x -3y = 13
4x - 3y = 11

Answer by checkley71(8403)

X-2Y=8
X+Y=-1 NOW SUBTRACT
------------------------------
-3Y=9
Y=9/-3
Y=-3 ANSWER.
X+(-3)=-1
X-3=-1
X=-1+3
X=2 ANSWER.
PROOF
2-2*-3=8
2+6=8
8=8